Question 3 : The heat of combustion for 1 mole of carbon to carbon dioxide is -410 kJ. How many kJ of heat would be liberated from the complete combustion of 60.0 g of carbon?
650.2 kJ
-2050 kJ
-24,600 kJ
-82 kJ

Question 4 : If the reaction A + B C has an equilibrium constant greater than one, which of the following statements is correct?
The reaction is product favoured.
The forward rate of reaction is fast.
The reaction is not spontaneous.
The backward rate of reaction is slow.
1
Expert's answer
2017-03-10T10:40:06-0500
Question 3: C + O2 = CO2 , ∆H = -410 kJ
n(C)=m/M=60/12=5 mole
1 mole – (-410 kJ)
5 mole – (x kJ)
x=(5∙(-410))/1=-2050 kJ
So the correct answer: -2050 kJ.
Question 4: If the reaction A + B = C is the equilibrium constant greater than one, it means that the concentration (amount of substance) products is greater than the concentration (amount of substance) reagents:
K_c=([C])/([A]∙[B])
So the correct answer: the reaction is product favoured.

Answer:
Question 3: -2050 kJ;
Question 4: the reaction is product favoured.